Benjamin has a 0.7 chance of passing his driving test the first time and a 0.85 chance of passing the second time. What is the probability of his passing on either the first or second try?

The probability of passing at either the first or second time round is P(passing first time) + P(taking the test for a second time then passing)
= 0.7 + (0.85 x 0.3)= 0.7 + 0.255= 0.955
Clarification: you must multiply the probability of Ben passing the second time by the probability that he failed the test the first time round. Then you can add this to the probability that he passes the first time.

RF

Related Maths GCSE answers

All answers ▸

Write down the coordinates of the turning point of the graph y = x^2 – 8x + 25


Find an expression for the nth term of this sequence: 3 - 11 - 19 - 27 - 35 . The nth term of a different sequence is 2n^3 + 3. Write down the first 3 terms of this sequence.


What is the lowest common multiple and the highest common factor of 120 and 150?


If x>2y and 2x<5y, what is the range of x when y=6?